# Localization of information driven by stochastic resetting

### v1

@@ Abstract
The dynamics of extended many-body systems are generically chaotic. Classically, a hallmark of chaos is the exponential sensitivity to initial conditions captured by positive Lyapunov exponents. Supplementing chaotic dynamics with stochastic resetting drives a sharp dynamical phase transition: we show that the Lyapunov spectrum, i.e., the complete set of Lyapunov exponents, abruptly collapses to zero above a critical resetting rate. At criticality, we find a sudden loss of analyticity of the velocity-dependent Lyapunov exponent, which we relate to the transition from ballistic scrambling of information to an arrested regime where information becomes exponentially localized over a characteristic length diverging at criticality with an exponent $\nu=1/2$. We illustrate our analytical results on generic chaotic dynamics by numerical simulations of coupled map lattices.
@@ Introduction.
At late times, and in the thermodynamic limit, we assume that the spatiotemporal growth of the OTOC is translation- and reflection-symmetric  [^1] If starting from an initial configuration that breaks translation or reflection symmetries, we assume that both symmetries are restored by the dynamics at late times, or that an appropriate ensemble average statistically restores them., and described by the ansatz (at $r=0$)
@@
for $|x-x^{\prime}|<v_{\rm LR}t$, while $D(x,x^{\prime};t)=0$ outside the causal light cone. Here, $\lambda(v)$ is the VDLE, first introduced as the co-moving [46, 47] or convective Lyapunov exponent [36]. This ansatz is known to capture the phenomenology of a wide class of classical and semi-classical chaotic systems [46, 48, 49, 50, 51]. Reflection symmetry implies that $\lambda(v)$ is an even function of $v$. We assume it is maximal at $v=0$ with $\lambda_{0}:=\lambda(0)>0$ coinciding with the largest Lyapunov exponent, and monotonically decreasing for $v>0$. In systems with local dynamics, $\lambda(v)$ is generally expected to be analytic around $v=0$. Therefore, we parametrize its small-velocity expansion as
@@
with $v_{0}>0$. The function $\lambda(v)$ vanishes at a finite butterfly velocity $v_{\rm B}:=\lambda^{-1}(0)$, which sets the maximum propagation speed of infinitesimal perturbations. The late-time dynamics feature a ballistic wavefront propagating at $v_{\rm B}$, with the exponential growth rate at the front given by $\lambda_{\rm B}=-v_{\rm B}\,\lambda^{\prime}(v_{\rm B})>0$. Except for the mild assumptions above, our results do not require further specifying the functional form of $\lambda(v)$. In App. C, we detail the special case of a parabolic VDLE of the form $\lambda(v)=\lambda_{0}[1-(v/v_{\rm B})^{2}]$ that can be treated analytically without any approximation. In Fig. 1, we numerically validate the ansatz in Eq. (4) and extract $\lambda(v)$ in the specific case of the CLM where the data are consistent with Eq. (5). See also Ref. [47].
@@
Noteworthy enough, for a discrete one-dimensional chain, where the Brillouin zone is restricted to $k\in[-\pi,\pi)$, one recovers the same expression as in Eq. (7), see App. B for details.
@@
At late times, we find that $\tilde{D}(x,x^{\prime};t)\sim\exp[\tilde{\lambda}(\frac{x-x^{\prime}}{t})t]$ where the functional form of the renormalized VDLE $\tilde{\lambda}(v)$ crucially depends on the relative distance to a critical resetting rate set by the largest Lyapunov exponent in the absence of resetting, $r_{\rm c}=\lambda_{0}$. The details are provided in App. A. For $r\leq r_{\rm c}$, the VDLE is simply shifted down by the resetting rate,
@@
In particular, the largest Lyapunov exponent under resetting $\lambda_{0}-r$ is positive: while the exponential sensitivity to initial conditions is reduced by a finite resetting rate, the dynamics are still chaotic. As the critical point is approached from below, $r\to r_{\rm c}^{-}$, the renormalized butterfly velocity vanishes as $\tilde{v}_{\rm B}\sim(r_{\rm c}-r)^{1/2}$. See Ref. [28] for a dynamical characterization of this chaotic regime. For $r>r_{\rm c}$, we find a more involved reduction of the VDLE, with
@@
Let us first remark that while averaging over resetting histories (or any other statistical ensemble) erases the information about individual trajectories, the notion of exponential sensitivity to initial conditions remains meaningful and can still be probed with OTOCs. This is notably the point of view taken in semiclassical approaches to quantum chaos. We go one step beyond and define the Lyapunov spectrum in the presence of resetting, $\tilde{\Lambda}(k)$, through Eq. (2) where $D$ is now replaced by $\tilde{D}$ defined in (8). We refer the reader to the App. B for the details of the computations. We find a spectral phase transition:
@@
$\xi_{r}:=-\frac{1}{\lambda^{\prime}(v^{*})}\stackrel{{\scriptstyle r\to r_{\rm c}^{+}}}{{\longrightarrow}}\frac{1}{2}\left(\frac{v_{0}^{2}}{\lambda_{0}}\frac{1}{r-r_{\rm c}}\right)^{1/2}.$ (17)
Equations (16) and (17) demonstrate that the spatiotemporal spread of information is brought to a complete halt in time, with a residual spatial smearing occurring on a length scale $\xi_{r}$ which diverges in approaching criticality from above with an exponent $\nu=1/2$. Note that Eq. (17) can be generalized to cases in which the small-velocity expansion in Eq. (5) starts at order $\alpha>2$, i.e. $\lambda(v\simeq 0)\simeq\lambda_{0}(1-|v/v_{0}|^{\alpha})$, yielding the critical $\nu=(\alpha-1)/\alpha$.
@@
At first sight, the information-localized regime for $r\geq r_{\rm c}$ resembles integrable dynamics: vanishing Lyapunov exponents lead to a collapsed spectrum, much like in systems constrained by an extensive number of conservation laws. Yet, the more compelling connection is with measurement-induced phase transitions (MIPTs) in quantum chaotic systems [5, 56, 57]. Classical constructions—via large-$N$ limits, inter-replica dynamics, or learnability transitions in monitored stochastic processes [50, 58, 59, 60]—have so far failed to reproduce the robust dynamical exponent $z=1$ signaling emergent space–time symmetry in MIPTs. Resetting dynamics, with its abrupt projection onto a restricted subset of configurations, naturally echoes features of the Born rule and can even generate long-range correlations [61, 62, 63, 64] reminiscent of quantum measurement protocols such as teleportation. This perspective suggests an interesting direction: probing the stability of the information-localized phase under partial resetting [65, 66], where only a fraction of the degrees of freedom are reset.
@@ Appendix A Velocity-dependent Lyapunov exponent under resetting
$\tilde{D}(x,x^{\prime};t)\sim\exp\left[\tilde{\lambda}\left(\frac{x-x^{\prime}}{t}\right)t\right],$ (20)
@@
$D(x,x^{\prime};t)\sim\exp\left[\lambda\left(\frac{x-x^{\prime}}{t}\right)t\right],$ (21)
where $\lambda(v)$ is assumed to be a continuous and even function of $v$ that monotonously decreases from $\lambda_{0}:=\lambda(v=0)>0$. Working at late times, let us postulate that we can use the ansatz (21) in the renewal equation given in Eq. (9) of the main text. This assumption will be checked a posteriori. Doing so, we obtain
$\tilde{D}(x;t)\sim r\int_{0}^{t}\!\!{\rm d}\tau\,{\rm e}^{[\lambda(x/\tau)-r]\tau}+{\rm e}^{[\lambda(x/t)-r]t}.$ (22)
@@
$\tilde{D}(v;t)\sim r\int_{0}^{t}\!\!{\rm d}\tau\,{\rm e}^{[\lambda(vt/\tau)-r]\tau}+{\rm e}^{[\lambda(v)-r]t}.$ (23)
@@
$0=\partial_{\tau}\left([\lambda(vt/\tau)-r]\tau\right).$ (24)
@@
$\lambda(v^{*})-r=v^{*}\lambda^{\prime}(v^{*}),$ (25)
Importantly, the existence of the saddle point is conditioned to $\tau^{*}<t$ or, equivalently, $v<v^{*}$. Otherwise, the integrand in Eq. (23) is simply maximal at its bound $\tau=t$ and we obtain $\tilde{D}(v>v^{*};t)\sim{\rm e}^{[\lambda(v)-r]t}$. A careful inspection of Eq. (25), using $\lambda^{\prime}(v)<0$, shows that the existence of a saddle point $v^{*}$ also requires $r\geq r_{\rm c}$ with the critical resetting rate $r_{\rm c}=\lambda_{0}$. In this case, the renormalized OTOC is approximated as the sum of two competing exponentials
$\tilde{D}(v<v^{*};t)\sim A\,r\,{\rm e}^{\lambda^{\prime}(v^{*})vt}+{\rm e}^{[\lambda(v)-r]t},$ (26)
@@
$\lambda^{\prime}(v^{*})\geq\frac{\lambda(v)-r}{v},$ (27)
the first term in Eq. (26) dominates. In the case $r\leq r_{\rm c}$, there is no saddle point and the integrand in Eq. (23) is simply maximal at its bound $\tau=t$ and $\tilde{D}(v;t)\sim{\rm e}^{[\lambda(v)-r]t}$. In all cases, the location of the maximum scales linearly with $t$. This justifies, a posteriori, the use of the bare velocity-dependent Lyapunov ansatz inside the renewal equation: the early-time behavior of the bare OTOC is irrelevant to the renormalized OTOC at late times. Altogether, restoring the possibility of $v<0$, we obtain the renomalized VDLE
$\tilde{\lambda}(v)=\begin{cases}\lambda(v)-r&\mbox{if }r\leq r_{\rm c}\\ \begin{cases}\lambda^{\prime}(v^{*})\,|v|&\mbox{for }|v|\leq v^{*}\\ \lambda(v)-r&\mbox{for }|v|\geq v^{*}\end{cases}&\mbox{if }r>r_{\rm c},\end{cases}$ (28)
where $v^{*}$ is determined by Eq. (25). Importantly, in the case $r>r_{\rm c}$, the renormalized VDLE develops linear branches for $|v|<v^{*}$. Note that $\tilde{\lambda}(v)$ is continuous but not analytic at $v=0$ and $v=\pm v^{*}$. As one approaches the critical resetting rate from above, the linear branch of the renormalized VDLE is controlled by the small-velocity expansion of $\lambda(v)$
$\lambda(v\simeq 0)=\lambda_{0}\,[1-(v/v_{0})^{2}]+O(v^{4}),$ (29)
@@
$v^{*}\stackrel{{\scriptstyle r\to r_{\rm c}^{+}}}{{\longrightarrow}}v_{0}\left(\frac{r-r_{\rm c}}{\lambda_{0}}\right)^{1/2}.$ (30)
@@
$\tilde{\lambda}_{0}=\max(\lambda_{0}-r,0).$ (31)
@@ Appendix B Lyapunov spectrum
$D(x,x^{\prime};t)\sim\exp\left[\lambda\left(\frac{x-x^{\prime}}{t}\right)t\right],$ (32)
@@
Noting that $D(x,x^{\prime};t)$ is a real symmetric square matrix in its spatial indices, this ensures that its eigenvalues are real. Furthermore, the translational invariance ensures that the standard momentum basis diagonalizes the matrix. Therefore, the Oseledets’ expression of the Lyapunov spectrum in Eq. (2) simplifies as
$\Lambda(k)=\lim\limits_{t\to\infty}\frac{1}{t}\log\left|\int{\rm d}x\,{\rm e}^{-{\rm i}kx}\exp\left[\lambda\left(\frac{x}{t}\right)t\right]\right|,$ (33)
@@
$\Lambda(k)=\lim\limits_{t\to\infty}\frac{1}{t}\log\left|\int{\rm d}v\,{\rm e}^{-{\rm i}kvt}\exp\left[\lambda\left(v\right)t\right]\right|.$ (34)
@@
$\lambda(v\simeq 0)=\lambda_{0}\,[1-(v/v_{0})^{2}]+O(v^{4}),$ (35)
@@
We now apply this formalism to the case of dynamics subject to stochastic resetting. Mutatis mutandis, the Lyapunov spectrum is expressed via the generalization of Eq. (34) as
$\tilde{\Lambda}(k)=\lim\limits_{t\to\infty}\frac{1}{t}\log\left|\int{\rm d}v\,{\rm e}^{-{\rm i}kvt}\exp\left[\tilde{\lambda}\left(v\right)t\right]\right|.$ (37)
where $\tilde{\lambda}(v)$ is the renormalized velocity-dependent Lyapunov computed in App. A.
In the case $r<r_{\rm c}=\lambda_{0}$, where $\tilde{\lambda}(v)=\lambda(v)-r$, Eq. (37) readily yields the renormalized spectrum
$\tilde{\Lambda}(k)=\Lambda(k)-r$ (38)
for all $k$, where $\Lambda(k)$ is the spectrum in the absence of stochastic resetting expressed in Eq. (34).
@@
$\tilde{\lambda}(v)=\lambda^{\prime}(v^{*})|v|\mbox{ for }v<v^{*},$ (39)
@@
$\int{\rm d}v\,{\rm e}^{-{\rm i}kvt}\exp\left[\lambda^{\prime}(v^{*})|v|t\right]=-\frac{2}{t}\frac{\lambda^{\prime}(v^{*})}{k^{2}+\lambda^{\prime}(v^{*})^{2}},$ (40)

### v2

@@ Abstract
The dynamics of extended many-body systems are generically chaotic. Classically, a hallmark of chaos is the exponential sensitivity to initial conditions captured by positive Lyapunov exponents. Supplementing chaotic dynamics with stochastic resetting drives a sharp dynamical phase transition: We show that the Lyapunov spectrum, i.e., the complete set of Lyapunov exponents, abruptly collapses to zero above a critical resetting rate. At criticality, we find a sudden loss of analyticity of the velocity-dependent Lyapunov exponent, which we relate to the transition from ballistic scrambling of information to an arrested regime where information becomes exponentially localized over a characteristic length diverging at criticality with an exponent $\nu=1/2$ and a dynamical exponent $z=2$. We illustrate our analytical results on generic chaotic dynamics by numerical simulations of coupled map lattices.
@@ Introduction.
At late times, and in the thermodynamic limit, we assume that the spatiotemporal growth of the OTOC is translation and reflection symmetric  [^1] If starting from an initial configuration that breaks translation or reflection symmetries, we assume that both symmetries are restored by the dynamics at late times, or that an appropriate ensemble average statistically restores them., and described by the ansatz (at $r=0$)
@@
for $|x-x^{\prime}|<v_{\rm LR}t$, while $D(x,x^{\prime};t)=0$ outside the causal light cone. Here, $\lambda(v)$ is the VDLE, first introduced as the comoving [46, 47] or convective Lyapunov exponent [36]. This ansatz is known to capture the phenomenology of a wide class of classical and semiclassical chaotic systems [46, 48, 49, 50, 51]. Reflection symmetry implies that $\lambda(v)$ is an even function of $v$. We assume it is maximal at $v=0$ with $\lambda_{0}:=\lambda(0)>0$ coinciding with the largest Lyapunov exponent, and monotonically decreasing for $v>0$. In systems with local dynamics, $\lambda(v)$ is generally expected to be analytic around $v=0$. Therefore, we parametrize its small-velocity expansion as
@@
with $v_{0}>0$. The function $\lambda(v)$ vanishes at a finite butterfly velocity $v_{\rm B}:=\lambda^{-1}(0)$, which sets the maximum propagation speed of infinitesimal perturbations. The late-time dynamics feature a ballistic wavefront propagating at $v_{\rm B}$, with the exponential growth rate at the front given by $\lambda_{\rm B}=-v_{\rm B}\,\lambda^{\prime}(v_{\rm B})>0$. Except for the mild assumptions above, our results do not require further specifying the functional form of $\lambda(v)$. In the App. C, we detail the special case of a parabolic VDLE of the form $\lambda(v)=\lambda_{0}[1-(v/v_{\rm B})^{2}]$ that can be treated analytically without any approximation. In Fig. 1, we numerically validate the ansatz in Eq. (4) and extract $\lambda(v)$ in the specific case of the CLM where the data are consistent with Eq. (5). See also Ref. [47].
@@
Noteworthy enough, for a discrete one-dimensional chain, where the Brillouin zone is restricted to $k\in[-\pi,\pi)$, one recovers the same expression as in Eq. (7), see App. D for details.
@@
At late times, we find that $\tilde{D}(x,x^{\prime};t)\sim\exp[\tilde{\lambda}(\frac{x-x^{\prime}}{t})t]$ where the functional form of the renormalized VDLE $\tilde{\lambda}(v)$ crucially depends on the relative distance to a critical resetting rate set by the largest Lyapunov exponent in the absence of resetting, $r_{\rm c}=\lambda_{0}$. The details are provided in App. A. Let us emphasize that $\tilde{\lambda}(v)$ does not directly relate to the standard definition of Lyapunov exponents since the average over trajectories is performed at the level of the OTOC rather than to its logarithm. While averaging over resetting histories (or any other statistical ensemble) erases the information about individual trajectories, the notion of exponential sensitivity to initial conditions remains meaningful and can still be probed with OTOCs. This is notably the point of view taken in semiclassical approaches to quantum chaos. For $r\leq r_{\rm c}$, the VDLE is shifted down by the resetting rate,
@@
The simplicity of this shift is rooted in the competing exponentials: Lyapunov growth and the Poisson process. In particular, the largest Lyapunov exponent under resetting $\lambda_{0}-r$ is positive: While the exponential sensitivity to initial conditions is reduced by a finite resetting rate, the dynamics are still chaotic. As the critical point is approached from below, $r\to r_{\rm c}^{-}$, the renormalized butterfly velocity vanishes as $\tilde{v}_{\rm B}\sim(r_{\rm c}-r)^{1/2}$. See Ref. [28] for a dynamical characterization of this chaotic regime. For $r>r_{\rm c}$, we find a more involved reduction of the VDLE, with
@@
We define the Lyapunov spectrum in the presence of resetting, $\tilde{\Lambda}(k)$, through Eq. (2) where $D$ is now replaced by $\tilde{D}$ defined in (8). We refer the reader to App. B for the details of the computations. We find a spectral phase transition,
@@
$\xi_{r}:=-\frac{1}{\lambda^{\prime}(v^{*})}\stackrel{{\scriptstyle r\to r_{\rm c}^{+}}}{{\longrightarrow}}\frac{1}{2}\frac{v_{0}}{\lambda_{0}}\left(\frac{\lambda_{0}}{r-r_{\rm c}}\right)^{1/2}.$ (17)
The approach to the steady-state is controlled by the timescale $\tau_{r}:=1/(r-r_{\rm c})$. Equations (16) and (17) demonstrate that the spatiotemporal spread of information is brought to a complete halt in time, with a residual spatial smearing occurring on a length scale $\xi_{r}$ which diverges in approaching criticality from above with an exponent $\nu=1/2$ and a dynamical exponent $z=2$. Note that Eq. (17) can be generalized to cases in which the small-velocity expansion in Eq. (5) starts at order $\alpha>2$, i.e. $\lambda(v\simeq 0)\simeq\lambda_{0}(1-|v/v_{0}|^{\alpha})$, yielding the critical $\nu=(\alpha-1)/\alpha$.
@@
At first sight, the information-localized regime for $r\geq r_{\rm c}$ resembles integrable dynamics: Vanishing Lyapunov exponents lead to a collapsed spectrum, much as in systems constrained by an extensive number of conservation laws. Yet, the more compelling connection is with measurement-induced phase transitions (MIPTs) in quantum chaotic systems [5, 56, 57, 58]. Classical constructions—via large-$N$ limits, inter-replica dynamics, or learnability transitions in monitored stochastic processes [50, 59, 60, 61]—have so far failed to reproduce the robust dynamical exponent $z=1$ signaling emergent spacetime symmetry in MIPTs. Resetting dynamics, with its abrupt projection onto a restricted subset of configurations, naturally echoes features of the Born rule and can even generate long-range correlations [62, 63, 64, 65] reminiscent of quantum measurement protocols such as teleportation. This perspective suggests an interesting direction: Probing the stability of the information-localized phase under partial resetting [66, 67], where only a fraction of the degrees of freedom are reset.
@@ Appendix A Velocity-dependent Lyapunov exponent under resetting
$\tilde{D}(x,x^{\prime};t)\sim\exp\left[\tilde{\lambda}\left(\frac{x-x^{\prime}}{t}\right)t\right],$ (A1)
@@
$D(x,x^{\prime};t)\sim\exp\left[\lambda\left(\frac{x-x^{\prime}}{t}\right)t\right],$ (A2)
where $\lambda(v)$ is assumed to be a continuous and even function of $v$ that monotonously decreases from $\lambda_{0}:=\lambda(v=0)>0$. Working at late times, let us postulate that we can use the ansatz (A2) in the renewal equation given in Eq. (9) of the main text. This assumption will be checked a posteriori. Doing so, we obtain
$\tilde{D}(x;t)\sim r\int_{0}^{t}\!\!{\rm d}\tau\,{\rm e}^{[\lambda(x/\tau)-r]\tau}+{\rm e}^{[\lambda(x/t)-r]t}.$ (A3)
@@
$\tilde{D}(v;t)\sim r\int_{0}^{t}\!\!{\rm d}\tau\,{\rm e}^{[\lambda(vt/\tau)-r]\tau}+{\rm e}^{[\lambda(v)-r]t}.$ (A4)
@@
$0=\partial_{\tau}\left([\lambda(vt/\tau)-r]\tau\right).$ (A5)
@@
$\lambda(v^{*})-r=v^{*}\lambda^{\prime}(v^{*}),$ (A6)
Importantly, the existence of the saddle point is conditioned to $\tau^{*}<t$ or, equivalently, $v<v^{*}$. Otherwise, the integrand in Eq. (A4) is simply maximal at its bound $\tau=t$ and we obtain $\tilde{D}(v>v^{*};t)\sim{\rm e}^{[\lambda(v)-r]t}$. A careful inspection of Eq. (A6), using $\lambda^{\prime}(v)<0$, shows that the existence of a saddle point $v^{*}$ also requires $r\geq r_{\rm c}$ with the critical resetting rate $r_{\rm c}=\lambda_{0}$. In this case, the renormalized OTOC is approximated as the sum of two competing exponentials
$\tilde{D}(v<v^{*};t)\sim A\,r\,{\rm e}^{\lambda^{\prime}(v^{*})vt}+{\rm e}^{[\lambda(v)-r]t},$ (A7)
@@
$\lambda^{\prime}(v^{*})\geq\frac{\lambda(v)-r}{v},$ (A8)
the first term in Eq. (A7) dominates. In the case $r\leq r_{\rm c}$, there is no saddle point and the integrand in Eq. (A4) is simply maximal at its bound $\tau=t$ and $\tilde{D}(v;t)\sim{\rm e}^{[\lambda(v)-r]t}$. In all cases, the location of the maximum scales linearly with $t$. This justifies, a posteriori, the use of the bare velocity-dependent Lyapunov ansatz inside the renewal equation: The early-time behavior of the bare OTOC is irrelevant to the renormalized OTOC at late times. Altogether, restoring the possibility of $v<0$, we obtain the renomalized VDLE
$\tilde{\lambda}(v)=\begin{cases}\lambda(v)-r&\mbox{if }r\leq r_{\rm c}\\ \begin{cases}\lambda^{\prime}(v^{*})\,|v|&\mbox{for }|v|\leq v^{*}\\ \lambda(v)-r&\mbox{for }|v|\geq v^{*}\end{cases}&\mbox{if }r>r_{\rm c},\end{cases}$ (A9)
where $v^{*}$ is determined by Eq. (A6). Importantly, in the case $r>r_{\rm c}$, the renormalized VDLE develops linear branches for $|v|<v^{*}$. Note that $\tilde{\lambda}(v)$ is continuous but not analytic at $v=0$ and $v=\pm v^{*}$. As one approaches the critical resetting rate from above, the linear branch of the renormalized VDLE is controlled by the small-velocity expansion of $\lambda(v)$
$\lambda(v\simeq 0)=\lambda_{0}\,[1-(v/v_{0})^{2}]+O(v^{4}),$ (A10)
@@
$v^{*}\stackrel{{\scriptstyle r\to r_{\rm c}^{+}}}{{\longrightarrow}}v_{0}\left(\frac{r-r_{\rm c}}{\lambda_{0}}\right)^{1/2}.$ (A11)
@@
$\tilde{\lambda}_{0}=\max(\lambda_{0}-r,0).$ (A12)
@@ Appendix B Lyapunov spectrum
$D(x,x^{\prime};t)\sim\exp\left[\lambda\left(\frac{x-x^{\prime}}{t}\right)t\right],$ (A13)
@@
Noting that $D(x,x^{\prime};t)$ is a real symmetric square matrix in its spatial indices, this ensures that its eigenvalues are real. Furthermore, the translational invariance ensures that the standard momentum basis diagonalizes the matrix. Therefore, the Oseledets’ expression of the Lyapunov spectrum in Eq. (2) of the main text simplifies as
$\Lambda(k)=\lim\limits_{t\to\infty}\frac{1}{t}\log\left|\int{\rm d}x\,{\rm e}^{-{\rm i}kx}\exp\left[\lambda\left(\frac{x}{t}\right)t\right]\right|,$ (A14)
@@
$\Lambda(k)=\lim\limits_{t\to\infty}\frac{1}{t}\log\left|\int{\rm d}v\,{\rm e}^{-{\rm i}kvt}\exp\left[\lambda\left(v\right)t\right]\right|.$ (A15)
@@
$\lambda(v\simeq 0)=\lambda_{0}\,[1-(v/v_{0})^{2}]+O(v^{4}),$ (A16)
@@
$\Lambda(k\simeq 0)$ $=\lambda_{0}-\lambda_{0}\left(\frac{v_{0}k}{2\lambda_{0}}\right)^{2}+\ldots,$ (A17)
where the ellipsis denotes higher-order terms in $k$ arising from subleading contributions in Eq. (A16).
@@
We now apply this formalism to the case of dynamics subject to stochastic resetting. Mutatis mutandis, the Lyapunov spectrum is expressed via the generalization of Eq. (A15) as
$\tilde{\Lambda}(k)=\lim\limits_{t\to\infty}\frac{1}{t}\log\left|\int{\rm d}v\,{\rm e}^{-{\rm i}kvt}\exp\left[\tilde{\lambda}\left(v\right)t\right]\right|,$ (A18)
where $\tilde{\lambda}(v)$ is the renormalized velocity-dependent Lyapunov computed in Sect. A.
In the case $r<r_{\rm c}=\lambda_{0}$, where $\tilde{\lambda}(v)=\lambda(v)-r$, Eq. (A18) readily yields the renormalized spectrum
$\tilde{\Lambda}(k)=\Lambda(k)-r$ (A19)
for all $k$, where $\Lambda(k)$ is the spectrum in the absence of stochastic resetting expressed in Eq. (A15).
@@
$\tilde{\lambda}(v)=\lambda^{\prime}(v^{*})|v|\mbox{ for }v<v^{*},$ (A20)
@@
$\int{\rm d}v\,{\rm e}^{-{\rm i}kvt}\exp\left[\lambda^{\prime}(v^{*})|v|t\right]=-\frac{2}{t}\frac{\lambda^{\prime}(v^{*})}{k^{2}+\lambda^{\prime}(v^{*})^{2}},$ (A21)